your family goes to a restaurant for dinner.There are 6 people in your family.Some order the chicken dinner for $14.80 and some order the steak dinner for $17.If the total bill was $91, how many people ordered each type of dinner?
Answer by richard1234(7193)

Note that the total bill is an integer amount. If we assume that c is the number of people who order a chicken dinner, we can say that c, 2c, 3c, 4c, nor 6c do not give us integer values for the total bill. Therefore there must have been five people who ordered the chicken bill. Checking,

5(14.80) + 17 = 74 + 17 = 91.